Technical Staff Blog

Category archives: Scheduled Maintance

Announcements about upcoming scheduled maintenance.

RSS feed of Scheduled Maintance

Last update on .

The CIS Networking Group will be swapping out a UPS in the 5th floor networking closet on Tuesday, Oct. 25th, 6am-7am. The networking switches in the closet will loose power during this maintenance.  These switches provide networking connections to southwest section of the CIT building on the 4th and 5th.

The following rooms on the ...

Last update on .

CIS Networking group is planning to make changes to the 10G switches, cit-cs-r14-sw-1 and cit-cs-r14-sw-2.  The switch config change requires a reboot.  They will be changing and rebooting the cit-cs-r14-sw-1 on Tuesday, June 21st at 5:30am and cit-cs-r14-sw-2 on Thursday, June 23rd at 6:00am.  The GPFS NSD nodes are connected to these switches ...

Last update on .

EMC_Isilon.png On Saturday June 4th, CIS will be physically moving their EMC Isilon, the hardware that provides file services for the campus. The move is scheduled to commence at 6am and complete around 4pm. We currently have three production services relying on the Isilon: email lists, ownCloud, and time machine backups. In order to keep our ...

Last update on .

network-2400px.png This was supposed to be a trouble free morning... just power up the switch we moved downstairs yesterday and add our redundancy back in. It seems to be a trend with this project, things not going quite as we planned...

07:14: The second distribution switch was powered on at 7am. Unfortuantely, something happened that ...

Last update on .

network-2400px.png On May 31st @ 7am CIS will be shutting down one of our two distribution switches and, because it is directly attached, we will be shutting down our backup firewall. Because of redundancy, we do not expect this work to cause any network outages. CIS will spend the remainder of May 31st moving the switch to ...

Last update on .

ibm-gpfs.jpg Overall, the migration of GPFS to the basement went remarkably well! We lost a total of two power supplies, both failures we could sustain through redundant hardware, and ran into a few bumps along the road, but it could have been a catastrophic failure and sure wasn't. We do have some additional cleanup that ...

Last update on .

ibm-gpfs.jpg The GPFS hardware is scheduled to be moved starting May 23rd @ 06:30am. Please check this page on the day of the move for periodic status updates...

... and so it begins...

05:45: Prep work for the move has begun. We are working to ensure that the most critical services remain active during the file ...

Last update on .

ibm-gpfs.jpg In just under two weeks, Tstaff will be physically moving our GPFS file system hardware from CIT 531 to the datacenter in the basement of the CIT. The purge of the upstairs machine rooms is the culmination of nearly two years of work by Tstaff. We will be shutting down GPFS @ 6:30am to begin ...

Last update on .

owncloud.png We are migrating the background data store for the ownCloud file syncing service today. The service will be offline while we perform a final sync. We estimate the final sync will take about an hour and a half, at which time we will enable the service again.

07:36: Taking the service offline for final ...

Last update on .

openvpn.png We downgraded one of our VPN servers to the previous version debian, in hopes that this version of OpenVPN doesn't suffer from the file handle leak bug. This morning, we failed our VPN service to this downgraded server. So far, it looks like this version is immune from the bug. We are going to ...